OpenStudy (anonymous):

Solve for J: 2/5 + J = 4/5.

OpenStudy (anonymous):

J = 6/5 J = 1/5 J = 2/5 J = -2/5 answers

OpenStudy (texaschic101):

2/5 + j = 4/5 j = 4/5 - 2/5 can you take it from here ?